# Symbolic Planning and Multi-Agent Path Finding in Extremely Dense Environments with Movable Obstacles

### v1

@@ 8 $P\leftarrow P\cup p_{i}$
4.3 PDDL-based Configuration Space Search
The configuration space search can also be described using PDDL. We define five predicates: (fre ?u), (tgt ?u), (blo ?u), (cmp ?u), and (goal ?u), indicating whether a vertex $u\in V$ contains a free space, target block, non-target block, completed target, or potential goal, respectively. The predicate (adjacent ?u ?v) denotes that vertices $u$ and $v$ are adjacent to each other, i.e., $(u,v)\in E$. Three actions are defined in the PDDL domain: (move_blo ?u ?v) and (move_tgt ?u ?v) for moving non-target and target blocks, respectively, and (complete ?u) for completion. The full PDDL domain is detailed in Appendix A.
In the problem file, vertices serve as objects, denoted as node-0-0, node-0-1, etc., excluding vertices containing obstacles. The initial state specifies each node’s type as either (tgt node-x-y), (blo node-x-y), or (fre node-x-y). All vertices except target blocks are marked with (cmp node-x-y), and valid terminal locations are marked with (goal node-x-y).
The goal state requires all vertices to be marked as (cmp node-x-y), with total action cost as the optimization metric. Note that this formulation does not limit the number of simultaneous moves. However, the fast-downward solver used in Sec. 5 creates a totally-ordered plan where only one action is applied a time.
@@ 4.4 Heuristic Approach: Moving Along the Least Blocking Path
Given the exponentially growing configuration search space with respect to the block number, algorithms in the previous three sections do not scale efficiently to large grids with numerous target blocks. In this section, we propose a pure heuristic-based approach that moves each target along the least-blocking vertex path.
Algorithm 2 outlines the heuristic approach. The process begins by assigning priorities to target blocks using the method described in Sec. 4.2. As illustrated in Figure 5, for each target, the planner first creates a least-blocking path, $U_{i}=\{u_{i}^{1},\cdots,u_{i}^{|U_{i}|}\}$, to the goal vertices. Then, for each vertex along this path, $u_{i}^{l}\in U_{i}$, the planner moves the closest blank to $u_{i}^{l}$ and subsequently moves the target block to the created blank, thus generating a block movement plan. A least-blocking path is defined as the lowest-cost path where target blocks, non-target blocks, and blanks are assigned different traversal costs.
During planning, the plan for the current target $i$ is created based on the map status where all previous target block plans have been executed. However, executing the plan for target $i$ does not require waiting for complete execution of all previous plans. Instead, it only requires waiting for the completion of plans affecting vertices that will be used by plan $p_{i}$. The function $t=\psi(u)$ takes a vertex $u\in V$ and returns this waiting time. By ensuring actions in $p_{i}$ occur no earlier than times specified by $\psi$, the planner enables parallel execution of multiple target movements.
@@
2Input: Grid $G(V,E)$, target set $\mathcal{I}$, goal vertices $V_{i}(i\in\mathcal{I})$
@@ 13 $P\leftarrow P\cup p_{i}\cup\textit{complete}{(i,t,u)}$
Figure 5: Move target blocks along the least blocking path.
@@ 4.5 LaCAM-based Multi-Agent Path Finding
By discretising time into timesteps and allowing multiple blocks to move to empty vertices simultaneously in each timestep, the problem is similar to a Multi-agent Path Finding problem. With the recent advances in solving large scale and high density MAPF problems, we introduce a Lazy Constraints Addition Search (LaCAM) (Okumura 2023b, a, 2024) based algorithm to solve the BRaP in this section.
@@
2Input: Grid $G(V,E)$, target set $\mathcal{I}$, non-target set $\mathcal{J}$, goal vertices $V_{i}$ $(i\in\mathcal{I})$, priorities $\omega$, current states $\pi$, next states $\pi^{\prime}$, current temp goals $g$,
@@
4 for $i\in\mathcal{I}$ do
5 if $\pi_{i}=g_{i}$ then $g_{i}\leftarrow\text{null}$
6  ; //Release temp goal
7 $g_{i}^{\prime}\leftarrow null$; //Next temp goal
8
@@ 7 $g_{i}^{\prime}\leftarrow null$; //Next temp goal
10for each block $k\in\mathcal{I}$ in priority order do
11 if $\pi_{k}^{\prime}=\perp$ then  $\textit{BRaP{}\_PIBT}(\text{null},k)$
12
13
14BRaP_PIBT(parent, block ii): begin
@@ 15 $\pi_{i}^{\prime}\leftarrow\pi_{i}$; //Default action wait
16
17 if $i\in\mathcal{I}$ then  $g_{i}^{\prime}\leftarrow$ if $g_{i}$ is null or occupied in $g_{i}^{\prime}$ then  $\text{closest\ goal\ }v$, $v\notin g^{\prime}$
@@ 31    $j\leftarrow$ the block occupies $u$ in $\pi$
34       if  $\pi^{\prime}_{j}\neq\perp$ or not $\textit{BRaP{}\_PIBT}(i,j)$ then continue
35
@@
42 return false
43
44
45return $\pi^{\prime}$, $g^{\prime}$
Algorithm 3 BRaP-PIBT
BRaP-PIBT Configuration Generator
@@
With the existence of the following conflict, PIBT (Okumura et al. 2022) is no longer applicable as a configuration generator for LaCAM in BRaP. This is because PIBT recursively decides the actions for chains of movable blocks, while, with following conflicts, only the blocks next to empty cells are movable. The additional complexity is the BRaP has multiple goal vertices for each target block, and the configuration generator needs to decide which goal vertex to move towards. In this section, we introduce BRaP-PIBT as the configuration generator for a LaCAM based MAPF solution for BRaP.
@@
The core intuition of the BRaP-PIBT algorithm is that every block takes turns to call for empty cells in a priority-based manner, ensuring that each block is processed exactly once per step to maintain algorithmic efficiency. This approach ensures that all blocks have the opportunity to make progress toward their goals by dynamically updating their priorities. Algorithm 3 describes the process of BRaP-PIBT.
@@
The BRaP-PIBT algorithm operates as a single-step planner that generates the next configuration state by iteratively assigning actions and temporary goals to blocks based on dynamically managed priorities. The algorithm maintains three key components: Priority Management, Recursive Planning and Temporary Goal Assignment.
Priority Management
@@
Similar to PIBT, each block $i$ maintains a priority value $\omega_{i}$ that determines the order in which blocks are processed. Priorities are dynamically updated based on goal achievement and movement status. When a block $i$ reaches any vertex in $V_{i}$, its priority is reset to a floating number in $(0,1)$. When the block $i$ is displaced from its goal vertices $V_{i}$, its priority is incremented by $1$. This priority update process happens at the beginning of each configuration generation phase (Algorithm 3, line 3).
Recursive Planning
@@
The algorithm processes target blocks according to their priority (Algorithm 3, lines 33). For each block, it attempts to move to the most preferred adjacent location. If the desired spot is occupied, the algorithm makes a recursive call on the blocking block, finding a chain of displacement requests that reaches an empty cell (Algorithm 3, lines 33). During this process, the algorithm returns the default action wait if no empty vertices are left or the chain of displacement is not possible. It then identifies available neighboring vertices and sorts them based on action preference metrics. The wait action is only considered as a preferred option if the block is not being actively pushed by another block to find the chain (Algorithm 3 line  3). The process ensures that a block remains stationary unless a move is strategically advantageous for itself or required to accommodate another block’s movement.
Temporary Goal Assignment
@@
For target blocks, BRaP-PIBT assigns temporary goals to guide them. When a block decides its action, a temporary goal is selected from the closest unassigned goal locations if the block do not have one or if its goal was taken away by a higher priority agent (Algorithm 3, line 3). The goal is released when a block reaches the assigned temporary goal (Algorithm 3, line 3).
@@
The action preference assignment differentiates between target blocks and non-target blocks. Target blocks prioritize moves that minimize the distance to their goals, while also considering proximity to empty spaces for tie breaking. Non-target blocks focus primarily on maintaining access to empty spaces to facilitate overall system mobility. The algorithm’s efficiency stems from its guarantee that each block is processed at most once per timestep. When a block is called recursively to facilitate movements, it is immediately assigned an action, thus is marked as processed.
@@
The BRaP-PIBT configuration generator integrates seamlessly with the LaCAM framework. For each high-level node expansion, LaCAM uses BRaP-PIBT to generate a single successor configuration state. If explored state are generated and low-level constraints are imposed, LaCAM forces BRaP-PIBT to satisfy the low-level constraints by first setting the next states in $\pi^{\prime}$ as the constrained locations of constrained blocks, then calling BRaP-PIBT for decisions on unconstrained blocks. The combination of LaCAM and BRaP-PIBT enables the rapid exploration of the configuration space while still maintaining LaCAM’s completeness guarantee.
@@ 5 Results and Discussion
Table 2: Experimental Setup Parameters
@@
Target number Minimum: 1
Maximum: 12.5% of grid vertices*
Blank number Minimum: 1
@@
Goal R1: Random goals, equal to # of targets
Goal R2: Random goals, twice the # of targets
@@
*For Goal B, the maximum target number is limited to either 12.5% of $|V|$ or twice the grid height, whichever is smaller.
Table 3: Algorithm Performance Metrics Across Different Configurations Grouped by Grid Sizes and Goal Types
Groups MAPF Heuristic Priority Config PDDL
@@
*Success rate, mean composite cost ratio, and mean makespan ratio of the proposed algorithms.The subscript values are standard deviations. Smaller ratios indicate better performance. Bold values indicate the best performance across all algorithms. Ratio metrics are omitted for success rates below 20%. The algorithm abbreviations are as follows. MAPF: LaCAM-based MAPF approach, Heuristic: heuristic approach, Priority: priority-based configuration space search, Config: configuration space search, PDDL: PDDL-based configuration space search.
@@
In this section, we evaluate the solution quality, computational time, and scalability of the algorithms proposed in Sec. 4 using an extensive set of 13,860 test cases. We systematically vary grid size, number of target blocks, number of blanks, and goal location selection methods. Table 2 lists the parameters used for generating these test cases, with representative examples visualized in Appendix D. Goal type B represents boundary vertex goals, while types R1 and R2 denote randomly selected goals, with type R1 being more restrictive in selection. Obstacles are modeled as a square region occupying 1/5 of the grid length, positioned in the bottom right corner of the grid. The PDDL formulation is solved using fast-downward [^1] Fast-downward: https://fast-downward.org. All algorithms are evaluated on a single machine with AMD EPYC 7R13 Processor. The time limit for one test case is 10 seconds.
@@
For each test case, success indicates finding a feasible plan within the time limit. To compare solution quality across algorithms, we establish a baseline for each test instance using the algorithm that finds the lowest final composite cost. Cost ratios are then calculated by dividing the final composite costs of other algorithms by this baseline (resulting in ratios always greater than 1). Makespan ratios are calculated in a similar way.
@@
Table 3 summarizes the success rate, composite cost ratio, and makespan ratio of the proposed algorithms across different grid configurations (13,860 test cases). The MAPF-based algorithm achieves the highest success rate, followed by the heuristic, priority, config, and PDDL algorithms. Only MAPF-based and heuristic algorithms successfully generate solutions for large instances. Overall, MAPF-based and heuristic algorithms achieve the lowest composite cost and makespan ratios, followed by priority, config, and PDDL approaches. Note that, the MAPF-based algorithm exhibits a higher standard deviation in composite cost on grid sizes $\times20$ to $\times80$ and for goal type R1. This variability indicates that for certain instances, its solutions can be substantially worse than its average performance.
@@
As shown in Table 3, algorithm success rates decrease with increasing grid size. Only the MAPF and heuristic algorithms successfully solve large instances (grid size $\geq$ $\times20$). It is important to note that the maximum number of targets increases proportionally with grid size. Notably, MAPF failures occur only when the grid size or target number exceeds computational capacity within the 10-second time limit, confirming its completeness as an algorithm - it fails only due to time constraints rather than algorithmic limitations. Table 3 also demonstrates that algorithms achieve higher success rates with goal types B and R2. Goal type R1, due to its more restrictive selection criteria, results in relatively lower success rates.
@@
Figure 6 illustrates the time required for each algorithm to report its first solution. The MAPF and heuristic-based algorithms find solutions within 1 millisecond for over 50% of the test cases. MAPF requires more than 1 second to create a solution in less than 10% of the cases.

### v2

@@ 4.4 Heuristic Approach: Moving Along the Least Blocking Path
Given the exponentially growing configuration search space with respect to the block number, algorithms in the previous three sections do not scale efficiently to large grids with numerous assigned blocks. In this section, we propose a pure heuristic-based approach that moves each assigned block along the least-blocking vertex path.
Algorithm 2 outlines the heuristic approach. The process begins by allocating priorities to assigned blocks using the method described in Sec. 4.3. As illustrated in Figure 5, for each assigned block, the planner first creates a least-blocking path, $U_{i}=\{u_{i}^{1},\cdots,u_{i}^{|U_{i}|}\}$, to the goal vertices. Then, for each vertex along this path, $u_{i}^{l}\in U_{i}$, the planner create an empty vertex at $u_{i}^{l}$ by iteratively moving blocks to the closest empty vertex and subsequently moves the assigned block to the emptied vertex $u_{i}^{l}$, thus generating a block movement plan. A least-blocking path is defined as the lowest-cost path where assigned blocks, unassigned blocks, and empty vertices are assigned different traversal costs.
During planning, the plan for the current assigned block $i$ is created based on the map status where all previous assigned block plans have been executed. However, executing the plan for assigned block $i$ does not require waiting for complete execution of all previous plans. Instead, it only requires waiting for the completion of plans affecting vertices that will be used by plan $p_{i}$. The function $t=\psi(u)$ takes a vertex $u\in V$ and returns this waiting time. By ensuring actions in $p_{i}$ occur no earlier than times specified by $\psi$, the planner enables parallel movements of multiple assigned block.
@@
2Input: Grid $G(V,E)$, assigned block set $\mathcal{I}$, goal vertices $V_{i}(i\in\mathcal{I})$
@@ 13 $P\leftarrow P\cup p_{i}\cup\textit{complete}{(i,t,u)}$
Figure 5: Move assigned blocks along the least blocking path.
@@ 4.5 LaCAM-based Multi-Agent Path Finding
By discretising time into timesteps and allowing multiple blocks to move to empty vertices simultaneously in each timestep, the problem is similar to a Multi-agent Path Finding problem. With the recent advances in solving large scale and high density MAPF problems, we introduce Block Rearrangement LaCAM, a Lazy Constraints Addition Search (LaCAM) (Okumura 2023b, a, 2024) based algorithm to solve the Block Rearrangement Problem in this section.
@@
2Input: Grid $G(V,E)$, assigned block set $\mathcal{I}$, unassigned set $\mathcal{J}$, goal vertices $V_{i}$ $(i\in\mathcal{I})$, priorities $\omega$, current states $\pi$, next states $\pi^{\prime}$, current temp goals $g$,
@@
4
[…]
6
7for each block $k\in\mathcal{I}$ in descending priority order do
8 if $\pi_{k}^{\prime}=\perp$ then $\textit{BR\_PIBT}(\text{null},k)$ ;
@@
10
11return $\pi^{\prime}$, $g^{\prime}$
12BR_PIBT(parent, block ii): begin
@@ 13 $\pi_{i}^{\prime}\leftarrow\pi_{i}$; //Default action wait
14
15 if $i\in\mathcal{I}$ then $g_{i}^{\prime}\leftarrow$ if $g_{i}$ is null or occupied in $g^{\prime}$ then $\text{closest\ goal\ }v\in V_{i}$, $v\notin g^{\prime}$
@@ 29    $j\leftarrow$ the block occupies $u$ in $\pi$
32       if $\pi^{\prime}_{j}\neq\perp$ or not $\textit{BR\_PIBT}(i,j)$ then continue
33
@@
42
43UpdatePriorities($\mathcal{I},V_{i},\omega,\pi$): begin
44 for each $i\in\mathcal{I}$ do
45    $\omega_{i}\leftarrow$ if $\pi_{i}\in V_{i}$ or $\omega_{i}=null$ then $rand(0,1)$
46    else $\omega_{i}+1.0$
47
48
49
50AssignPreferences($u,i,g^{\prime}_{i}$): begin
51 $u.preference\leftarrow$ if $i\in\mathcal{I}$ then $(Distance(u,g_{i}^{\prime}),DistanceToEmpty(u))$
52 else $(DistanceToEmpty(u),random())$
53
54
Algorithm 3 BR-PIBT
BR-PIBT Configuration Generator
@@ 45    $\omega_{i}\leftarrow$ if $\pi_{i}\in V_{i}$ or $\omega_{i}=null$ then $rand(0,1)$
With the existence of following conflicts, PIBT (Okumura et al. 2022) is no longer applicable as a configuration generator for LaCAM in BRaP. This is because PIBT recursively decides the actions for chains of movable blocks, while, due to following conflicts, only the blocks next to empty vertices are movable. The additional complexity is the BRaP has multiple goal vertices for each assigned block, and the configuration generator needs to decide which goal vertex to move towards. In this section, we introduce BR-PIBT as the configuration generator for Block Rearrangement LaCAM (BR-LaCAM).
@@
The core intuition of the BR-PIBT algorithm is that every block takes turns to call for empty vertices in a priority-based manner, ensuring that each block is processed exactly once per step to maintain algorithmic efficiency. This approach ensures that all blocks have the opportunity to make progress toward their goals by dynamically updating their priorities. Algorithm 3 describes the process of BR-PIBT.
@@
The BR-PIBT algorithm operates as a single-step planner that generates the next configuration state by iteratively assigning actions and temporary goals to blocks based on dynamically managed priorities. The algorithm maintains three key components: 1) Priority Management, 2) Recursive Planning and 3) Temporary Goal Allocation.
@@
1) Priority Management: Similar to PIBT, each block $i$ maintains a priority value $\omega_{i}$ that determines the order in which blocks are processed. Priorities are dynamically updated based on goal achievement and movement status. When a block $i$ reaches any vertex in $V_{i}$, its priority is reset to a floating number in $(0,1)$. When the block $i$ is displaced from its goal vertices $V_{i}$, its priority is incremented by $1$. This priority update process happens at the beginning of each configuration generation phase (Algorithm 3 line 3 and 3).
@@
2) Recursive Planning: The algorithm processes assigned blocks according to their priority (Algorithm 3 lines 33). For each block, it attempts to move to the most preferred adjacent location. If the desired spot is occupied, the algorithm makes a recursive call on the blocking block, finding a chain of displacement requests that reaches an empty vertex (Algorithm 3, lines 33). During this process, the algorithm returns the default action wait if no empty vertices are left or the chain of displacement is not possible. It then identifies available neighboring vertices and sorts them based on action preference metrics (Algorithm 3 lines 33). The wait action is only considered as a preferred option if the block is not being actively pushed by another block to find the chain (Algorithm 3 line  3). The process ensures that a block remains stationary unless a move is strategically advantageous for itself or required to accommodate another block’s movement.
@@
3) Temporary Goal Allocation: For assigned blocks, BR-PIBT allocates temporary goals to guide them. When a block decides its action, a temporary goal is selected from the closest unallocated goal locations if the block does not have one or if its goal was taken away by a higher priority block (Algorithm 3 line 3).
@@
The action preference differentiates between assigned blocks and unassigned blocks (Algorithm 3 line 3). Assigned blocks prioritize moves that minimize the distance to their goals, while also considering proximity to empty vertices for tie breaking. Unassigned blocks focus primarily on maintaining access to empty vertices to facilitate overall system mobility. The algorithm’s efficiency stems from its guarantee that each block is processed at most once per timestep. When a block is called recursively to facilitate movements, it is immediately decided an action, thus is marked as processed.
@@
The BR-PIBT configuration generator integrates seamlessly with the LaCAM framework. For each high-level node expansion, LaCAM uses BR-PIBT to generate a single successor configuration state. If explored state are generated and low-level constraints are imposed, LaCAM forces BR-PIBT to satisfy the low-level constraints by first setting the next states in $\pi^{\prime}$ as the constrained locations of constrained blocks, then calling BR-PIBT for decisions on unconstrained blocks. The combination of LaCAM and BR-PIBT, called BR-LaCAM, enables the rapid exploration of the configuration space while still maintaining LaCAM’s completeness guarantee.
Claim 1: BR-LaCAM is complete, which returns a solution for solvable Block Rearrangement Problems.
Proof.
@@ 5 Results and Discussion
[…]
(g) $\times10$, 13 assigned,
3 empties
(h) $\times10$, 1 assigned,
1 empty vertex
(i) $\times10$, 13 assigned,
8 empties
(j) $\times20$, 40 assigned, 10 empties
(k) $\times40$,
160 assigned, 40 empties
(l) $\times80$, 128 assigned,
160 empties
(m) $\times80$, 128 assigned,
1600 empties
(n) $\times80$, 800 assigned,
160 empties
(o) $\times80$, 800 assigned,
1600 empties
Figure 6: Example test cases. Dark gray vertices indicate the goals for the assigned blocks. Black vertices indicate obstacles.
@@
Assigned blocks number Minimum: 1
Maximum: 12.5% of grid vertices
Empty vertex number Minimum: 1
@@
[…]
@@
Table 2: Experimental setup parameters. For Goal B, the maximum assigned block number is limited to either 12.5% of $|V|$ or twice the grid height, whichever is smaller.
Groups BR-LaCAM Heuristic Priority Config PDDL
@@
Table 3: Algorithm performance across different configurations grouped by grid sizes and goal types, including success rate, mean composite cost ratio, and mean makespan ratio. The subscript values are standard deviations. Smaller ratios indicate better performance. Bold values indicate the best performance across all algorithms. Ratio metrics are omitted for success rates below 20%. The algorithm abbreviations are as follows. BR-LaCAM: LaCAM with BR-PIBT, Heuristic: heuristic approach, Priority: priority-based configuration space search, Config: configuration space search, PDDL: PDDL-based configuration space search.
@@
In this section, we evaluate the solution quality, computational time, and scalability of the algorithms proposed in Sec. 4 using an extensive set of 13,860 test cases. We systematically vary grid size, number of assigned blocks, number of empty vertices, and goal location selection methods. Table 2 lists the parameters used for generating these test cases, with representative examples visualized in Figure 6. Goal type B represents boundary vertex goals, while types R1 and R2 denote randomly selected goals, with type R1 being more restrictive in selection. Obstacles are modeled as a square region occupying 1/5 of the grid length, positioned in the bottom right corner of the grid. The PDDL formulation is solved using fast-downward [^1] Fast-downward: https://fast-downward.org. All algorithms are evaluated on a single machine with AMD EPYC 7R13 Processor. The time limit for one test case is 10 seconds, in line with the typical convergence time of the algorithms.
@@
For each test case, success indicates finding a feasible plan within the time limit. To compare solution quality across algorithms, we establish a baseline for each test instance using the algorithm that finds the lowest final composite cost. Each algorithm’s final composite cost is divided by the baseline to calculate cost ratios (resulting in ratios always greater than 1). Makespan ratios are calculated in a similar way.
@@
Table 3 summarizes the success rate, composite cost ratio, and makespan ratio of the proposed algorithms across different grid configurations (13,860 test cases). The BR-LaCAM achieves the highest success rate, followed by the heuristic, priority, config, and PDDL algorithms. Only BR-LaCAM and heuristic algorithms successfully generate solutions for large instances. Overall, BR-LaCAM and heuristic algorithms achieve the lowest composite cost and makespan ratios, followed by priority, config, and PDDL approaches. Note that, the BR-LaCAM algorithm exhibits a higher standard deviation in composite cost on grid sizes $\times20$ to $\times80$ and for goal type R1. This variability indicates that for instances with goal vertices sparsely distributed and on large maps, BR-PIBT’s myopic behavior may leads to substantially higher solution costs than its average performance.
@@
As shown in Table 3, algorithm success rates decrease with increasing grid size. Only the BR-LaCAM and heuristic algorithms successfully solve large instances (grid size $\geq$ $\times20$). It is important to note that the maximum number of assigned blocks increases proportionally with grid size. Notably, BR-LaCAM failures occur only when the grid size or assigned block number exceeds computational capacity within the 10-second time limit, confirming its completeness as an algorithm - it fails only due to time constraints rather than algorithmic limitations. Not reflected in the table, increasing the number of empty vertices reduces search depth, empirically improving success rates and reducing both composite cost and makespan. Table 3 also demonstrates that algorithms achieve higher success rates with goal types B and R2. Goal type R1, due to its more restrictive selection criteria, results in relatively lower success rates. Figure 7 visualizes the composite cost ratios across all 13,860 test instances, with each instance represented by a single dot. Consistent with the results in Table 3, failure rates increase with grid size and number of assigned blocks. The Cost Ratios among algorithms typically range between 1 and 10 relative to the best performer.
Figure 7: Composite Cost Ratio for each instance, grouped by number of assigned blocks (AB) and grid size (GS). The number and dots in the no solution region shows how many instances are unsolved for each algorithm in each group.
@@
Figure 8 illustrates the time required for each algorithm to report its first solution. The BR-LaCAM and heuristic-based algorithms find solutions within 1 millisecond for over 50% of the test cases. BR-LaCAM’s initial solution time exceeds 1 second in fewer than 10% of cases.